# You Only Forward Once: An Efficient Compositional Judging Paradigm

### v2

@@ 4.2 Data Construction
Training set. To construct the training dataset, we use the SA-1B dataset [kirillov2023segment], introduced alongside the Segment Anything Model. It contains approximately 11M diverse, high-resolution images sourced from a privacy-preserving data engine and annotated with 1.1B high-quality segmentation masks. Since we do not require the masks, as shown in Fig. 4(a), we use only the images, prompting an MLLM to analyze them, propose properties that hold or do not hold for the corresponding images, and provide reasons. The detailed prompt is shown in Fig. 5(a). We randomly sample 1.2M images to ensure diversity and generalizability. For each image $\mathcal{I}_{i}$, the final sample is represented as $(\mathcal{I}_{i},(\mathcal{P}_{ij},\mathcal{A}_{ij},\mathcal{R}_{ij})_{j=1}^{N})$, denoting the image, the proposed properties, the binary answers(“yes” or “no”), and the reasons. To maintain a balanced and randomized mix of satisfied and unsatisfied properties, we prompt the MLLM to produce approximately equal numbers of each and shuffle them during training. Additionally, although we instruct the model to generate exactly ten properties per sample in the training set, the number of properties per training sample varies, typically remaining close to ten.. To evaluate the model and mitigate overfitting, we split the dataset at a ratio of $6:1$ into training and validation sets.
Test set. We use the LRVS-Fashion dataset [lepage2023lrvs], a large-scale public benchmark for referred visual search (RVS) in the fashion domain, to construct a test set. It comprises 272K fashion products and 842K images extracted from real-world catalogs. To evaluate generalization in Image-Text reranking, as depicted as Fig. 4(b), we construct image pairs and queries; the model then judges which of the two images better matches the corresponding query. Specifically, we randomly sample image pairs $(\mathcal{I}_{i1},\mathcal{I}_{i2})$ from the same category and, using an MLLM, generate a natural-language query $\mathcal{Q}_{i}$ that mimics how a customer would express their preferences. The MLLM then produces the answer $\mathcal{A}_{i}$ indicating whether the first image $\mathcal{I}_{i1}$ better matches $\mathcal{Q}_{i}$, the set of query requirements $(\mathcal{R}_{ij})_{j=1}^{M_{i}}$ (where $M_{i}$ is the number of requirements), and a valid python expression $\mathcal{E}_{i}$ used to compute the final recommendation score based on our model’s predictions. The detailed prompt is demonstrated in Fig. 5(b).
@@ 4.3 Evaluation Metrics
To evaluate the model’s ability to make judgements based on its previous ones, one more metrics is needed on dependency-aware judgement task, named Dependency-Aware Accuracy ($Acc_{\mathrm{dep}}$).
On the test sets, we evaluate whether the model can rank the image pairs properly. Therefore, we report the error rate of ranking $Error_{\mathrm{rank}}$, denoting the percentage of image pairs ranked wrongly.
@@ 4.4 Comparison on LAION-RVS-Fashion
In this section, we present an analysis of YOFO against state-of-the-art multimodal rerankers [jina2025blog].
LAION-RVS-Fashion is a specialized fashion dataset, which inherently exhibits a significant domain shift relative to our training set constructed entirely on the broad and diverse SA-1B corpus. Despite this distributional gap, YOFO demonstrates remarkable robustness and adaptability when evaluated on fashion-specific recommendation tasks. As presented in Tab. 2, YOFO not only achieves strong performance but consistently outperforms the strong baseline jina-reranker-m0, even though it has never been exposed to fashion-related data during training. This result highlights a key advantage of our approach: YOFO learns general-purpose judging capabilities that transfer effectively across domains. Crucially, it can be deployed directly in specialized subdomains—such as fashion—without any fine-tuning or domain adaptation. This zero-shot generalization underscores the model’s practical utility in real-world scenarios where labeled data is scarce or domain shifts are common, positioning YOFO as a versatile solution for cross-domain recommendation systems.
YOFO also delivers substantially higher throughput than generative rerankers. This efficiency derives from YOFO’s single-forward-pass design: unlike generative or multi-stage rerankers [ma2024fine, nogueira2019multi] that must process each requirement sequentially or in multiple rounds, YOFO produces all judgments in a single forward pass. Consequently, its inference latency increases only mildly with the number of requirements compared with generative or multi-stage rerankers, making it especially well suited to real-time, large-scale recommendation scenarios where both responsiveness and expressiveness are critical.
Additionally, the YOFO-trained Qwen3-VL-2B-Instruct model attains the lowest ranking error rate—1.1% lower than the model trained with Qwen2-VL-2B-Instruct—and achieves higher throughput, which we attribute to Qwen3-VL-2B’s smaller parameter count.
Methods $Error_{\mathrm{rank}}$ (%) $\downarrow$ $\mathrm{Throughput}$ (pairs/s) $\uparrow$
Jina-Reranker-m0 16.2 36.41
YOFO (Qwen2-VL) 4.8 35.08
YOFO (Qwen3-VL) 3.7 47.6
Table 2: Ranking Error Rates on LAION-RVS-Fashion252 Dataset. YOFO outperforms jina-reranker-m0, while maintaining high throughput.
@@ 4.5 Dependency-Aware Judgement
To evaluate whether our model can make judgments conditioned on previous ones, we randomly select two properties from each sample in the validation set and replace the second property with “The answer to this question is the opposite of the answer to the previous question”. The model is therefore required to infer the answer to the second question without being given the first property’s answer, while making both judgments simultaneously. We fine-tune the pretrained Qwen2-VL-2B-Instruct model on this dataset for one epoch, keep other settings unchanged, and then compare it with the model in Sec. 4.4. Although challenging, Tab. 3 demonstrates YOFO’s dependency-aware judgment ability under the dependent-question setup, approaching 100%. Without training on the purposely-designed dataset, YOFO cannot judge correctly, because the properties in the training and validation sets can be judged individually, rendering dependency-aware analysis unnecessary. The base model performs poorly on this task, as it has not been exposed to scenarios where previous tokens are unobserved when predicting the next token.
@@ 4.6 Ablation Study
Effectiveness of YOFO training. We fine-tune Qwen2-VL-2B-Instruct and Qwen3-VL-2B-Instruct models on the training set using LoRA [hu2022lora] for one epoch and evaluate it on the validation set. As shown in Tab. 4, after training, Property-wise Accuracy improves by over 20%, reaching above 90%, while Sample-wise Accuracy increases by approximately 40%. These ablation results demonstrate the effectiveness of our training paradigm. Moreover, the YOFO model trained with Qwen3-VL-2B-Instruct performs slightly better than its counterpart trained with Qwen2-VL-2B-Instruct. The Sample-wise Accuracy remains moderate because the training and validation sets contain, on average, nearly ten properties per sample, making it challenging to judge all properties correctly.
@@
Effectiveness of post-hoc CoT. As shown in the last row of Tab. 4, YOFO benefits from post-hoc CoT. This likely arises because the model is trained to generate and consider reasons before making judgments. This component is particularly useful in complex settings where detailed reasons can be annotated.
Rank of LoRA. We evaluate different ranks of LoRA, and the results in Tab. 5 show that $r=16$ and $r=64$ perform the best. To increase the model’s representational capacity, we adopt $r=64$ as our default experimental setting.
@@
(a) Image Pair and Query
(b) Jina-Reranker-M0’s Prediction
(c) YOFO’s Prediction
@@ 4.7 Additional Analysis
We present a case study in Fig. 6. The user seeks a sleeveless, asymmetric dress that is not striped and preferably has ruching. However, jina-reranker-m0 assigns a higher score to the dress in the second image, overlooking its stripes and lack of ruching. In contrast, YOFO judges, for each image, whether the query’s requirements are satisfied and correctly assigns a higher score to the first image. This case explicitly demonstrates the effectiveness and interpretability of our method.
We also analyze judgment accuracy across positions on the SA-1B validation set. Specifically, $Acc_{i}$ denotes the accuracy for $i$-th requirement, and Fig. 7 visualizes the results. The figure shows that the base model achieves over 80% accuracy on the first requirement, but its accuracy declines rapidly toward 70% thereafter. In contrast, YOFO maintains accuracy above 90%, with or without post-hoc CoT. This likely occurs because the base model’s behavior on the first requirement resembles autoregressive generation, yet it fails to adapt to our template-conditioned judgment paradigm in which the ground-truth labels for previous judgments are unobserved. After training, YOFO adapts to this paradigm and consequently maintains high accuracy across positions.

### v3

@@ 4.2 Data Construction
Training Set. To construct the training dataset, we use the SA-1B dataset (Kirillov et al., 2023), introduced alongside the Segment Anything Model. It contains approximately 11M diverse, high-resolution images sourced from a privacy-preserving data engine and annotated with 1.1B high-quality segmentation masks. Since we do not require the masks, as shown in Figure 4 (a), we use only the images, prompting an MLLM to analyze them, propose properties that hold or do not hold for the corresponding images, and provide reasons. We randomly sample 1.2M images to ensure diversity and generalizability. For each image $\mathcal{I}_{i}$, the final sample is represented as $(\mathcal{I}_{i},(\mathcal{P}_{ij},\mathcal{A}_{ij},\mathcal{R}_{ij})_{j=1}^{N})$, denoting the image, the proposed properties, the binary answers(“yes” or “no”), and the reasons. To maintain a balanced and randomized mix of satisfied and unsatisfied properties, we prompt the MLLM to produce approximately equal numbers of each and shuffle them during training. Additionally, although we instruct the model to generate exactly ten properties per sample in the training set, the number of properties per training sample varies, typically remaining close to ten.. To evaluate the model and mitigate overfitting, we split the dataset at a ratio of $6:1$ into training and validation sets.
Test Set. We use the LRVS-Fashion dataset (Lepage et al., 2023), a large-scale public benchmark for referred visual search (RVS) in the fashion domain, to construct a test set. It comprises 272K fashion products and 842K images extracted from real-world catalogs. To evaluate generalization in Image-Text reranking, as depicted as Figure 4 (b), we construct image pairs and queries; the model then judges which of the two images better matches the corresponding query. Specifically, we randomly sample image pairs $(\mathcal{I}_{i1},\mathcal{I}_{i2})$ from the same category and, using an MLLM, generate a natural-language query $\mathcal{Q}_{i}$ that mimics how a customer would express their preferences. The MLLM then produces the answer $\mathcal{A}_{i}$ indicating whether the first image $\mathcal{I}_{i1}$ better matches $\mathcal{Q}_{i}$, the set of query requirements $(\mathcal{R}_{ij})_{j=1}^{M_{i}}$ (where $M_{i}$ is the number of requirements), and a valid python expression $\mathcal{E}_{i}$ used to compute the final recommendation score based on our model’s predictions.
Data Curation. To ensure the quality of the training and test sets generated by MLLMs, we implement a human-in-the-loop curation protocol. For training set and test set independently, we iteratively refine the corresponding prompt, generate samples from a small, randomly selected subset of the source data, and manually evaluate sample quality and correctness against predefined standard. This iteration continues until the generated samples consistently satisfy the quality thresholds. The finalized prompts are then employed to produce the complete training and test sets, respectively.
@@ 4.3 Evaluation Metrics
To evaluate the model’s ability to make judgments conditioned on its previous decisions, we introduce an additional metric for the dependency-aware judgment task, termed Dependency-Aware Accuracy and denoted as $Acc_{\mathrm{dep}}$. It measures the percentage of questions that depend on previous ones and are answered correctly.
On the test sets, we evaluate whether the model can rank the image pairs properly. Therefore, we report the error rate of ranking $Err_{\mathrm{rank}}$, denoting the percentage of image pairs ranked wrongly.
@@ 4.4 Comparison on LAION-RVS-Fashion
In this section, we present a comparative analysis of YOFO with representative state-of-the-art embedding models (e.g., CLIP (Radford et al., 2021), SigLIP 2 (Tschannen et al., 2025) and BLIP-2 (Li et al., 2023)) and multimodal rerankers (e.g., Jina-Reranker-M0 (Jina AI, 2025), Qwen3-VL-Reranker (Li et al., 2026) and LamRA-Rank (Liu et al., 2025)). Notably, all evaluated rerankers are built on Qwen2-VL (Wang et al., 2024) or Qwen3-VL (Bai et al., 2025a).
LAION-RVS-Fashion is a specialized fashion dataset, which inherently exhibits a significant domain shift relative to our training set constructed entirely on the broad and diverse SA-1B corpus. Despite this distributional gap, YOFO demonstrates remarkable robustness and adaptability when evaluated on fashion-specific recommendation tasks. As shown in Table 1, YOFO achieves competitive performance and consistently outperforms all strong baselines—despite never being trained on fashion-related data. Notably, LamRA-Rank-7B (fine-tuned on Qwen2-VL-7B), despite its significantly larger parameter count, exhibits a ranking error rate nearly twice that of YOFO finetuned on Qwen2-VL-2B. Similarly, Qwen3-VL-Reranker-2B outperforms LamRA-Rank-7B, but still underperforms YOFO. This performance gap is attributable to the inherent limitation of single-score prediction paradigm in capturing fine-grained semantic relationships.
This result highlights a key advantage of our approach: YOFO learns general-purpose judging capabilities that transfer effectively across domains. Crucially, it can be deployed directly in specialized subdomains—such as fashion—without any fine-tuning or domain adaptation. This zero-shot generalization underscores the model’s practical utility in real-world scenarios where labeled data is scarce or domain shifts are common, positioning YOFO as a versatile solution for cross-domain recommendation systems.
Regarding throughput evaluation, throughput metrics for embedding models are omitted, as their embedding process is typically performed offline to accelerate retrieval in modern recommendation systems. Among the evaluated multimodal rerankers, YOFO finetuned on Qwen3-VL-2B achieves the highest throughput, demonstrating superior computational efficiency.
Additionally, YOFO based on Qwen3-VL-2B achieves a 1.1% lower ranking error rate and higher throughput compared to its Qwen2-VL-2B counterpart, which we attribute to Qwen3-VL-2B’s smaller parameter count.
Table 1: Ranking Error Rates on LAION-RVS-Fashion252 Dataset. YOFO achieves the lowest ranking error rate across all baseline models while simultaneously attaining the highest throughput among all evaluated rerankers.
Methods $Err_{\mathrm{rank}}$ (%) $\downarrow$ $\mathrm{Throughput}$ (pairs/s) $\uparrow$
Clip-ViT-L 42.0 -
Siglip2-ViT-G 48.3 -
Blip2-itm-ViT-G 37.4 -
Jina-Reranker-m0 16.2 36
Qw3-Reranker-2B 8.7 48
LamRA-Rank-7B 9.3 5
YOFO (Qwen2-VL) 4.8 35
YOFO (Qwen3-VL) 3.7 48
@@ 4.5 Dependency-Aware Judgement
To evaluate whether our model can make judgments conditioned on previous ones, we randomly select two properties from each sample in the validation set and replace the second property with “The answer to this question is the opposite of the answer to the previous question”. The model is therefore required to infer the answer to the second question without being given the first property’s answer, while making both judgments simultaneously. We fine-tune the pretrained Qwen2-VL-2B-Instruct model on this dataset for one epoch, keep other settings unchanged, and then compare it with the model in Section 4.4. Although challenging, Table 2 demonstrates YOFO’s dependency-aware judgment ability under the dependent-question setup, approaching 100%. Without training on the purposely-designed dataset, YOFO cannot judge correctly, because the properties in the training and validation sets can be judged individually, rendering dependency-aware analysis unnecessary. The base model performs poorly on this task, as it has not been exposed to scenarios where previous tokens are unobserved when predicting the next token.
@@ 4.6 Ablation Study
Effectiveness of YOFO training. We fine-tune Qwen2-VL-2B-Instruct and Qwen3-VL-2B-Instruct models on the training set using LoRA (Hu et al., 2022) for one epoch and evaluate it on the validation set. As shown in Table 3, after training, Property-wise Accuracy improves by over 20%, reaching above 90%, while Sample-wise Accuracy increases by approximately 40%. These ablation results demonstrate the effectiveness of our training paradigm. Moreover, the YOFO model trained with Qwen3-VL-2B-Instruct performs slightly better than its counterpart trained with Qwen2-VL-2B-Instruct. The Sample-wise Accuracy remains moderate because the training and validation sets contain, on average, nearly ten properties per sample, making it challenging to judge all properties correctly.
@@
Effectiveness of post-hoc CoT. As shown in the last row of Table 3, YOFO benefits from post-hoc CoT. This likely arises because the model is trained to generate and consider reasons before making judgments. This component is particularly useful in complex settings where detailed reasons can be annotated.
Rank of LoRA. We evaluate different ranks of LoRA, and the results in Table 4 show that $r=64$ perform the best. Therefore, we adopt $r=64$ as our default experimental setting.
@@ 4.7 Additional Analysis
We present a case study in Figure 5. The user seeks a sleeveless, asymmetric dress that is not striped and preferably has ruching. However, jina-reranker-m0 assigns a higher score to the dress in the second image, overlooking its stripes and lack of ruching. In contrast, YOFO judges, for each image, whether the query’s requirements are satisfied and correctly assigns a higher score to the first image. This case explicitly demonstrates the effectiveness and interpretability of our method.
We also analyze judgment accuracy across positions on the SA-1B validation set. Specifically, $Acc_{i}$ denotes the accuracy for $i$-th requirement, and Figure 6 visualizes the results. The figure shows that the base model achieves over 80% accuracy on the first requirement, but its accuracy declines rapidly toward 70% thereafter. In contrast, YOFO maintains accuracy above 90%, with or without post-hoc CoT. This likely occurs because the base model’s behavior on the first requirement resembles autoregressive generation, yet it fails to adapt to our template-conditioned judgment paradigm in which the ground-truth labels for previous judgments are unobserved. After training, YOFO adapts to this paradigm and consequently maintains high accuracy across positions.